a man of mass 60kg is standing on a boat of mas 140 kg, which is at rest in still water.the man is initially 20 m away from the shore. he starts walking on the boat for 4s with constant speed 1.5 m/s  towards the shore? what will be his final distance from the shore?

Grade:11

1 Answers

Sai Pavan
9 years ago
The  man moves with a speed of 1.5 m/sec for 4 sec . So distance travelled by the man during this time period is 6 m. Let us assume the shore is to our left , when the moves towards the shore there will be a relative displacement of boat towards right . So the relative displacement of boat is = md/(M+m) , where m is mass of man ,and d is the distance he walks towards the shore and M is the mass of boat .after substituting the values given we get the relative displacement to be1.8m.so the total distance from the shore is = initial distance from the shore – distance walked by the man + relative displacement of boat = 20-6+1.8=15.8m